OpenStudy (aizhalee):

ALGEBRA HELP PLEASE f(x)=3x, g(x) = x-5 My answer (f•g)(x)= 3(x-5) 3x-15 Am I right ?

OpenStudy (aizhalee):

@welshfella can you verify for me please if Im doing it right ?

OpenStudy (aizhalee):

@Mehek14

TheSmartOne (thesmartone):

hi

OpenStudy (aizhalee):

hello :) thank you stopping by to help

OpenStudy (aizhalee):

so that means [(3x)][(x-5)]

TheSmartOne (thesmartone):

my bad, you're correct :)

OpenStudy (aizhalee):

so my answer is correct? :)

TheSmartOne (thesmartone):

yes :)

TheSmartOne (thesmartone):

( f o g)(x) = f (g(x))

OpenStudy (aizhalee):

Yay ! Thanks a lot

TheSmartOne (thesmartone):

wait a seconf o.o

TheSmartOne (thesmartone):

second*

OpenStudy (aizhalee):

ok ;o

TheSmartOne (thesmartone):

http://prntscr.com/90zp62

TheSmartOne (thesmartone):

so we would have to do what I posted earlier :)

TheSmartOne (thesmartone):

3x(x-5) Just use the distributive laws a(b-c) = ab - ac

OpenStudy (aizhalee):

ok

OpenStudy (aizhalee):

can you explain on how that should be done ?

TheSmartOne (thesmartone):

The multiplication for 3x(x-5) ?

OpenStudy (aizhalee):

3x - 15?

TheSmartOne (thesmartone):

no, the x here are not multiplication (3x)(x-5) = (3x)(x) - (3x)(5) = ?

OpenStudy (aizhalee):

I know 3x * 5 = 15x . (3x)(x) = 3x

TheSmartOne (thesmartone):

(3x)(x) = 3x^2

OpenStudy (aizhalee):

Yess thank you .

TheSmartOne (thesmartone):

x = x^1 (x)(x) = (x^1)(x^1) = x^{1+1} = x^2

OpenStudy (aizhalee):

so to the 2nd power ?

TheSmartOne (thesmartone):

Yup

TheSmartOne (thesmartone):

Your final answer is 3x^2 - 15x

OpenStudy (aizhalee):

wow ! Ok Thank you so much for explaining I will be adding these to my notes

TheSmartOne (thesmartone):

Anytime! :)